OpenStudy (anonymous):

measure of angle B is 14 degrees more than 3x the measure of angle A. Measure of angle C is 46 degrees more than the angle of A. Find the measures of each angle.

OpenStudy (anonymous):

In an isosceles triangle, all 3 angles must add up to 180 degrees. angle B will be 180-46-14. It's really just that simple.

OpenStudy (anonymous):

so you would than subtract and that would be the answer right?

OpenStudy (anonymous):

that would make the answers b3 134, 120,

OpenStudy (anonymous):

B=3X+14 A=3X C=46+3X....SO.... 3X+14+3X+46+3X=180...9X+60=180...SUBTRACT 80 FROM BOTH SIDES...YOU GET 9X=120....SOLVE FOR X....X=13.333333...PLUG THAT INTO YOUR A,B,AND C VALUES AND YOU WILL GET B=54 A=40 AND C=86